OpenStudy (anonymous):

the sum of a 2-digit number and the 2-digit number when the digits are reversed are 77 and the difference is 45. what are the 2 two digit numbers

Parth (parthkohli):

10x + y + 10y + x = 77 10x + y - 10y - x = 45 Adding, 20x + 2y = 122 2(10x + y) = 122 10x + y = 61 Can you do it now?